A mathematics teacher in senior high school develop a problem-solving test to randomly selected 40 students. These students had an average score of 85 and standard deviation of 5. If this population had a mean score of 90 and standard deviation of 3, use 5% Level of Significance to test the hypothesis.


1
Expert's answer
2021-12-06T16:14:58-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu=90"

"H_1:\mu\not=90"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{85-90}{3/\sqrt{40}}\approx-10.541"

Since it is observed that "|z| = 10.541 > 1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=2P(z<-10.541)\approx0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 90, at the "\alpha = 0.05" significance level.